OpenStudy (anonymous):

Help answering (2x+10)/(2x-5)=4

OpenStudy (anonymous):

You can multiply both sides of the equation by 2x-5 (with the stipulation that x cannot be 5/2). That should give you something you can solve pretty quickly.

OpenStudy (anonymous):

u are in the university or school? because there are 2 ways of solving it :S

OpenStudy (anonymous):

college

OpenStudy (anonymous):

ok the u might have to pass the 4 to le left so it can be -4, the get de MCM

OpenStudy (anonymous):

u know the critic points?

OpenStudy (anonymous):

i think it's more to do with multiplying both sides so far i get, (2x+10)/(2x-5)=4 , so that divide (2x-5) both sides = 2x+10 = 6x + 10?

OpenStudy (anonymous):

4(2x-5) = 8x -20 not 6x + 10.

OpenStudy (anonymous):

So you will have 2x + 10 = 8x - 20

OpenStudy (anonymous):

so then 10+20 = 8x - 2x?

OpenStudy (anonymous):

which = 30 = 6x?

OpenStudy (anonymous):

and then x = 30/6?

OpenStudy (anonymous):

its ok but i was taught that if u dont know if x is negative u cannot do that :S thats why i asked u

OpenStudy (anonymous):

that right so far pol?

OpenStudy (anonymous):

Looks reasonable yes.

OpenStudy (anonymous):

so then x = 5 (i think)

OpenStudy (anonymous):

Indeed. And if you plug 5 back in for x in the original equation?

OpenStudy (anonymous):

\[{2(5) + 10\over 2(5) - 5} = 4 \implies {4(5) \over 5} = 4 \implies 4= 4\]

OpenStudy (anonymous):

Looks good.

OpenStudy (anonymous):

okay thanks for that :) ill post if need anymore help later on :)

OpenStudy (anonymous):

Certainly!